# On the numerical approximation of hyperbolic mean curvature flows for surfaces

### v1

@@ 4.2.1 Radial evolutions for a sphere
We begin with a convergence experiment in order to validate our schemes (3.1) and (3.16) with $g(s)=1+\tfrac{1}{2}s$, using the solution discussed in [18, §7.1]. In fact, a family of spheres in ${\mathbb{R}}^{3}$ with radius $r(t)$ and outer normal $\nu(t)$ is a solution to (1.3) with $\mathcal{V}_{\mathcal{S}\mid_{t=0}}=V_{0}\in{\mathbb{R}}$ if
@@
Thus, if $V_{0}>0$, then $r(t)$ will at first increase until it hits a maximum at $t=\tfrac{1}{2}r_{0}V_{0}$ after which it will decrease and shrink to a point at time $T_{max}$, where $T_{max}=\frac{r_{0}}{2}\bigl{(}V_{0}+\sqrt{V_{0}^{2}+2}\bigr{)}$, see [18, p. 611]. On the other hand, if $V_{0}\leq 0$ then the sphere will monotonically shrink to a point at time $T_{max}$.
@@
As discussed above, for an initial sphere, depending on the sign of $V_{0}\in{\mathbb{R}}$, the family of spheres either expands at first and then shrinks, or shrinks immediately. We visualize these different behaviours in Figure 9. In each case we observe a smooth solution until the spheres shrink to a point, meaning that $|\mathcal{S}^{m}_{h}|$ and $1/{\mathcal{K}}^{m}_{\infty}$ approach zero at the same time. Compared to Figure 2 we note small differences in time scales, and in the maximal radius reached for the simulation with $V_{0}=1$. We also note the highly oscillatory plots of $1/{\mathcal{K}}^{m}_{\infty}$ for nonzero $V_{0}$, which we believe to be numerical artifacts.
Figure 9: The flow (1.3) starting from a unit sphere. Above we show the evolutions of $|\mathcal{S}^{m}_{h}|$ (solid) and $A^{m}$ (dashed) over time for $V_{0}=0$, $V_{0}=1$ and $V_{0}=-1$. The final times for these computations are $T=0.69$, $T=1.35$ and $T=0.35$, respectively. Below we show the corresponding evolutions of $1/{\mathcal{K}}^{m}_{\infty}$ over time.
@@ 4.2.2 Axisymmetric evolutions for nonspherical surfaces
In analogy to the simulation in Figure 3, we begin with an experiment for an initial $2:1:1$ cigar, when the initial velocity is chosen as $V_{0}=0$, see Figure 10. Interestingly, the ensuing evolution is markedly different. In fact, there seems to be the onset of a singularity at the much earlier time $t=0.35$. Moreover, the part of the surface that now becomes flat is vertical, compared to the horizontal part that flattened in Figure 3.
Figure 10: The flow (1.3), with $V_{0}=0$, starting from a $2:1:1$ cigar. Above we visualize $\mathcal{S}^{m}_{h}$ from (3.1) at times $t=0,0.2,T=0.34$ in two ways: a frontal view and a cut through the $x_{1}$-$x_{2}$ plane. Below we show $x^{m}$ from (3.16) with $g(s)=1+\tfrac{1}{2}s$ at times $t=0,0.2,T$, as well as a plot with $|\mathcal{S}^{m}_{h}|$ (solid) and $A^{m}$ (dashed), and a plot of $1/\mathcal{K}^{m}_{\infty}$ (right) over time.

### v2

@@ 4.2.1 Radial evolutions for a sphere
We begin with a convergence experiment in order to validate our schemes (3.1) and (3.17) with $g(s)=1+\tfrac{1}{2}s$, using the solution discussed in [23, §7.1]. In fact, a family of spheres in ${\mathbb{R}}^{3}$ with radius $r(t)$ and outer normal $\nu(t)$ is a solution to (1.3) with $\mathcal{V}_{\mathcal{S}\mid_{t=0}}=V_{0}\in{\mathbb{R}}$ if
@@
Thus, if $V_{0}>0$, then $r(t)$ will at first increase until it hits a maximum at $t=\tfrac{1}{2}r_{0}V_{0}$ after which it will decrease and shrink to a point at time $T_{max}$, where $T_{max}=\frac{r_{0}}{2}\bigl{(}V_{0}+\sqrt{V_{0}^{2}+2}\bigr{)}$, see [23, p. 611]. On the other hand, if $V_{0}\leq 0$ then the sphere will monotonically shrink to a point at time $T_{max}$.
@@
To further investigate the suboptimal convergence rates, we repeat the experiments for Tables 3 and 4 but now choose $\Delta t=0.5h^{2}_{0}$ and $\Delta t=h^{2}$, respectively. The results are shown in Tables 5 and 6, demonstrating quadratic convergence rates. Hence, as expected, these results appear to confirm that the temporal discretization, and in particular the explicit approximation of $g(\mathcal{V}_{\mathcal{S}}^{2})$, is responsible for the reduced convergence rates seen in Tables 3 and 4.
$V_{0}=0$ $V_{0}=1$ $V_{0}=-1$
1536 2.0854e-01 4.6490e-03 1.62 7.2549e-03 2.12 1.4806e-02
6144 1.0472e-01 1.5653e-03 1.58 2.1150e-03 1.79 3.3757e-03 2.15
24576 5.2416e-02 4.4500e-04 1.82 6.3384e-04 1.74 8.1195e-04 2.06
98304 2.6215e-02 1.2257e-04 1.86 1.8075e-04 1.81 2.2052e-04 1.88
393216 1.3108e-02 3.3678e-05 1.86 5.0157e-05 1.85 5.9463e-05 1.89
Table 5: Same as Table 3 but with $\Delta t=0.5h^{2}_{0}$.
$V_{0}=0$ $V_{0}=1$ $V_{0}=-1$
32 4.1126e-04 5.8329e-05 4.4130e-04
64 1.0181e-04 2.01 6.9665e-06 3.07 1.1113e-04 1.99
128 2.5403e-05 2.00 1.1422e-06 2.61 2.6938e-05 2.04
256 6.3444e-06 2.00 2.1542e-07 2.41 6.7589e-06 1.99
512 1.5813e-06 2.00 4.2844e-08 2.33 1.6773e-06 2.01
Table 6: Same as Table 4 but with $\Delta t=h^{2}$.
As discussed above, for an initial sphere, depending on the sign of $V_{0}\in{\mathbb{R}}$, the family of spheres either expands at first and then shrinks, or shrinks immediately. We visualize these different behaviours in Figure 9. In each case we observe a smooth solution until the spheres shrink to a point, meaning that $|\mathcal{S}^{m}_{h}|$ and $1/{\mathcal{K}}^{m}_{\infty}$ approach zero at the same time. Compared to Figure 2 we note small differences in the time scales, and in the maximal radius reached for the simulation with $V_{0}=1$.
Figure 9: The flow (1.3) starting from a unit sphere. Above we show the evolutions of $|\mathcal{S}^{m}_{h}|$ (solid) and $A^{m}$ (dashed) over time for $V_{0}=0$, $V_{0}=1$ and $V_{0}=-1$. The final times for these computations are $T=0.7$, $T=1.36$ and $T=0.36$, respectively. Below we show the corresponding evolutions of $1/{\mathcal{K}}^{m}_{\infty}$ over time.
@@ 4.2.2 Axisymmetric evolutions for nonspherical surfaces
In analogy to the simulation in Figure 3, we begin with an experiment for an initial $2:1:1$ cigar, when the initial velocity is chosen as $V_{0}=0$, see Figure 10. Interestingly, the ensuing evolution is slightly different. Apart from the shorter time scale, the final shape of the surface appears to be nonconvex.
Figure 10: The flow (1.3), with $V_{0}=0$, starting from a $2:1:1$ cigar. Above we visualize $\mathcal{S}^{m}_{h}$ from (3.1) at times $t=0,0.2,T=0.39$ in two ways: a frontal view and a cut through the $x_{1}$-$x_{2}$ plane. Below we show $x^{m}$ from (3.17) with $g(s)=1+\tfrac{1}{2}s$ at times $t=0,0.2,T$, as well as a plot with $|\mathcal{S}^{m}_{h}|$ (solid) and $A^{m}$ (dashed), and a plot of $1/\mathcal{K}^{m}_{\infty}$ (right) over time.